OpenStudy (anonymous):

which sum is equal to x^2 +6x-5/ x^2-25

OpenStudy (anonymous):

@geishaology

OpenStudy (anonymous):

@Best_Mathematician help please

OpenStudy (anonymous):

Kay.

OpenStudy (anonymous):

factor numrator and denominator

OpenStudy (anonymous):

Is this a multiple choice or work problem?

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

That. . .did not answer my question lmao.

OpenStudy (anonymous):

@Best_Mathematician the denominator would be (x-5)(x+5)

OpenStudy (anonymous):

I don't know the rest

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

@Best_Mathematician

OpenStudy (anonymous):

@hba @.Sam. @vivianrush

OpenStudy (anonymous):

so what do u say guys

OpenStudy (anonymous):

answer is \[\frac{ 1 }{ x+5 } +\frac{ 5 }{ x-5 }+1\]

OpenStudy (anonymous):

@Best_Mathematician I don't have that in my answers

OpenStudy (anonymous):

or it should be \[\frac{ x(x+6)-5 }{ (x+5)(x-5) }\]

OpenStudy (anonymous):

@Best_Mathematician can u please check the choices and tell me which one is it cause that's also non of the choices

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@amistre64 I need ur help here plz

OpenStudy (amistre64):

it looks like they want you to decompose a fraction ....

OpenStudy (anonymous):

I think so

OpenStudy (amistre64):

have we factored them yet?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

these old eyes cant locate it in the post, can you show me so that i can see if its factored correctly?

OpenStudy (anonymous):

the denominator would be (x-5)(x+5)

OpenStudy (anonymous):

I don't know the numerator

OpenStudy (anonymous):

it's a multiable choice question

OpenStudy (amistre64):

i think you have a typo in the numerator

OpenStudy (anonymous):

sorry I didn't get it

OpenStudy (amistre64):

can you retype the numerator ... i think there might be a mistake in the one you posted

OpenStudy (amistre64):

not the options, the problem itself

OpenStudy (amistre64):

ok, then we can proceed

OpenStudy (anonymous):

this is the equation

OpenStudy (anonymous):

so which answer do u think

OpenStudy (amistre64):

lets split it up and try to recombine it \[\frac{x^2-6x+5}{(x-5)(x+5)}=\frac{A}{(x-5)}+\frac{B}{(x+5)}\] normally we would cross multiply A to B to "solve" \[x^2-6x+5=A(x+5)+B(x-5)\] \[x^2-6x+5=Ax+5A+Bx-5B\] the problem im seeing is that there is no x^2 on the right to compare with

OpenStudy (amistre64):

if we simply go with the ropes on this let x=5 \[\frac{5^2-6(5)+5}{\cancel{(x-5)(x+5)}}=\cancel{\frac{A}{(x-5)}}+\frac{B}{(5+5)}\] \[B=10(25-30+5)=0\] let x=-5 \[\frac{(-5)^2-6(-5)+5}{\cancel{(x-5)(x+5)}}=\frac{A}{(-5-5)}+\cancel{\frac{B}{(5+5)}}\] \[A=-10(25+30+5)=-600\]

OpenStudy (amistre64):

and those aint options, so im thinking the problem is simply posted bad to begin with in you material

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%28x%5E2-6x%2B5%29%2F%28x%5E2-25%29 the wolf agrees